OpenStudy (anonymous):

*Multiple Choice! Just need to make sure I'm right! Will give medal to first reply!* Dan bought a truck for $29,800. The value of the truck depreciated at a constant rate per year. The table below shows the value of the truck after the first and second years: Year: 1 2 Value (in dollars): 26,522 23,604.58

OpenStudy (anonymous):

Which function best represents the value of the truck after t years? A) f(t) = 29,800(0.89)t B) f(t) = 26,522(0.89)t C) f(t) = 29,800(0.11)t D) f(t) = 26,522(0.11)t

OpenStudy (anonymous):

I think it's the first one

OpenStudy (amistre64):

can you explain why you think its A?

OpenStudy (anonymous):

Because B and D don't have the beginning amount of 29,800

OpenStudy (amistre64):

good thinking :) so it has to be A or C. Assuming the the t is an exponent ... what is leaning you tpwards A?

OpenStudy (amistre64):

10% is 2980, which is not the value of the truck the first year, so it has to be the 90% is what im thinking as a logical solution

OpenStudy (anonymous):

So that makes it A?